# IBNorm: Information-Bottleneck Inspired Normalization for Representation Learning

### v1

@@ C.2 Results on ImageNet
C.3 The effect of Compression Operation in IBNorm
We instantiate three representative forms of the compression function $f_{\lambda}$: linear (IBNorm-S), logarithmic (IBNorm-L), and hyperbolic tangent (IBNorm-T).
@@ C.3 The effect of Compression Operation in IBNorm
These choices capture complementary compression behaviors. Specifically, IB-S provides a linear baseline, IB-L yields progressively stronger suppression of large activations, and IB-T enforces a hard saturation effect on extreme values. Together, they span the spectrum from mild to strong tail suppression while sharing the desirable properties of boundedness and monotonicity. While other monotonic compression functions are possible, we provide these three forms to demonstrate the range of behaviors and provide a principled comparison.
@@
In conclusion, for these functionals, the compression ratio $\alpha_{\lambda}$ in Eqn. (7) is given by $\alpha_{\lambda}=1/\lambda$, which guarantees the bounded compression property whenever $\lambda\geq 1$.
Compression operations in IBNorm compress the tail of activations while adjusting higher-order statistics. IBNorm-L, IBNorm-T, and IBNorm-S exhibit different abilities to compress the tails of activations.
@@
C.4 Estimated Mutual Information Quantities and Token-level IB Value across Training
We track the evolution of three mutual information quantities during training: predictive information $\hat{I}(Y;T_{l})$, task–nuisance information $\hat{I}(T_{l-1};T_{l})$, and their token-level IB value $\hat{I}(Y;T_{l})-\hat{I}(T_{l-1};T_{l})$. These quantities are measured throughout training on test dataset for Llama-130M and GPT-2 (124M) on C4 and OpenWebText, respectively. The details in token-level IB value calculation are in Appendix G.
@@ C.4 Estimated Mutual Information Quantities and Token-level IB Value across Training
Figure 6: Evolution of predictive information $\hat{I}(Y;T_{l})$, task–nuisance information $\hat{I}(T_{l-1};T_{l})$, and the token-level IB value during training of Llama-130M on C4, using LN, IBNorm-T, and NormalNorm. Results are reported on the test set.
@@
Figure 7: Evolution of predictive information $\hat{I}(Y;T_{l})$, task–nuisance information $\hat{I}(T_{l-1};T_{l})$, and the token-level IB value during training of GPT-2 (124M) on OpenWebText, using LN, IBNorm-L, and NormalNorm. Results are reported on the test set.
@@
As shown in Figs. 67, and 8, these results demonstrate that IBNorm encourages representations that retain task-relevant information while suppressing irrelevant task-nuisance information, providing a explanation for the observed empirical gains.
@@ Appendix D Details of Other Terms
In addition, we have the optimum–dependent term:
@@ E.1 Proof for Theorem 1
[…]
@@
Given the dataset $S\sim\mathbb{P}_{X,Y}$ with size $M$ and one-layer neural networks, $f_{\text{1}}=\Phi_{\text{1}}\circ h$. We analyze the IB value of intermediate features $\tilde{T}_{1}:=\psi\circ\zeta\circ h(X)$. First, we define the empirical gap $\Delta_{S}(\tilde{T}_{1};\tilde{T}_{1}^{*})$ between the IB bottleneck derived from arbitrary representation $\tilde{T}_{1}$ and the optimum representation $\tilde{T}_{1}^{*}$, i.e.
@@
The detailed proof is given in Appendix E.4.
Since $G(\cdot)$ is monotone in its variance arguments, Lemma 5 implies that, for a fixed trade-off hyperparameter $\beta\in[0,1]$ and confidence parameter $\delta\in(0,1)$, with probability at least $1-\delta$ over the sample set $S\sim\mathbb{P}(X,Y)$ with size $M$, the following inequality holds:
@@
[…]
@@ E.2 Proof for Lemma 3
Note that $\phi$ is a continuous, and concave function, and that $\lim\limits_{x\rightarrow 0}\phi(x)=0$. In addition, we have the following properties for the auxiliary function $\phi(\cdot)$.
@@
For terms A and B, we have the following bound which is given by the Theorem 2 in the paper (Shamir et al., 2010).
@@
$\quad+2\int_{\mathcal{T}}\phi\Bigg((2+\sqrt{2\log((|\mathcal{Y}|+2)/\delta)})\sqrt{\frac{V(p(T|X))}{M}}\Bigg)p_{T}(t)\,dt\quad(\text{Law of Large Numbers})$
@@
$\quad+2\int_{\mathcal{T*}}\phi\Bigg((2+\sqrt{2\log((|\mathcal{Y}|+2)/\delta)})\sqrt{\frac{V(p(T^{*}|X))}{M}}\Bigg)p_{T^{*}}(t)\,dt\quad(\text{Law of Large Numbers})$
@@
Thus, in our setting, we have:
@@
For terms A’ and B’, we have the following bound which is given by the Theorem 2 in the paper (Shamir et al., 2010).
@@
$|I(X;T)-\hat{I}(X;T)|\leq(2+\sqrt{2\log((|\mathcal{Y}|+2)/\delta)})\sqrt{\frac{V({H}(T|X))}{M}}$
@@
$|I(X;T^{*})-\hat{I}(X;T^{*})|\leq(2+\sqrt{2\log((|\mathcal{Y}|+2)/\delta)})\sqrt{\frac{V({H}(T^{*}|X))}{M}}$
@@
Thus, in our setting, we have
@@
Combining I and II, we have the bound
@@
where $Z(\tilde{T}_{0}^{*},\beta)$ is the partition function. Thus, we can write the left terms to optimum–dependent components:
@@ E.3 Proof for Proposition 4
$H(s_{\lambda}(Z))=H(Z)+\mathbb{E}[\log|\det J_{s_{\lambda}}(Z)|]\leq H(Z),$
because $|\det J_{s_{\lambda}}(Z)|=\prod_{i=1}^{d}f_{\lambda}^{\prime}(|Z_{i}-\mu|)\leq\prod_{i=1}^{d}\alpha_{\lambda}\leq 1$ by the bounded compression property of $f_{\lambda}$.
@@ E.4 Proof for Lemma 5
$\tilde{T}_{s}=\Phi_{\text{s}}(X^{\prime})=\eta\circ\psi_{\text{s}}\circ\zeta(X^{\prime}),\qquad\tilde{T}_{IB}=\Phi_{\text{IB}}(X^{\prime})=\eta\circ\psi_{\text{IB}}\circ\zeta(X^{\prime}),$
where $X^{\prime}=h(X)$, $\psi_{\text{IB}}=\psi_{\text{s}}\circ s_{\lambda}$ with $s_{\lambda}$ being the compression operation in IBNorm.
@@
$H(\tilde{T}_{IB}|X)\leq H(\tilde{T}_{s}|X),\qquad H(\tilde{T}_{IB}|Y)\leq H(\tilde{T}_{s}|Y).$
If we consider the entropy $H$ in the discrete case, we can directly derive the relationship: consider the conditional entropy as a random variable over $X$ (or $Y$). Let $f^{\prime}(Z):=H(s_{\lambda}(Z))$, then:
1. $s_{\lambda}$ is a Lipschitz map with Lipschitz constant $L\leq 1$ because $|\det J_{s_{\lambda}}(Z)|=\prod_{i=1}^{d}f_{\lambda}^{\prime}(|Z_{i}-\mu|)\leq\prod_{i=1}^{d}\alpha_{\lambda}\leq 1$ by the bounded compression property of $f_{\lambda}$.
2. $H(\cdot)$ is 1-Lipschitz continuous on the space of probability distributions bounded away from zero.
@@
In the differential entropy case, we cannot guarantee a global Lipschitz constant $L<1$ for $f^{\prime}(\tilde{T}_{s}|X)=H(\tilde{T}_{s}|X)+\Delta^{\prime}(\tilde{T}_{s}|X)$. Instead, we proceed in a high-probability setting.
@@
$p_{\tilde{T}_{IB}|X=x}(t)=p_{\tilde{T}_{s}|X=x}(s_{\lambda}^{-1}(t))\,|\det J_{s_{\lambda}^{-1}}(t)|,$
@@
$H(\tilde{T}_{IB}|X=x)=H(\tilde{T}_{s}|X=x)+\Delta(\tilde{T}_{s}),\quad\Delta(\tilde{T}_{s}):=\log|\det J_{s_{\lambda}}(\tilde{T}_{s})|.$
@@
$\mathbb{V}(H(\tilde{T}_{IB}|X)\mid\tilde{T}_{s}\in\mathcal{T}_{\delta}^{s})=\mathbb{V}(H(\tilde{T}_{s}|X)+\Delta(\tilde{T}_{s})\mid\tilde{T}_{s}\in\mathcal{T}_{\delta}^{s}).$
Expanding using the variance formula,
$\mathbb{V}(H(\tilde{T}_{s}|X)+\Delta(\tilde{T}_{s}))=\mathbb{V}(H(\tilde{T}_{s}|X))+\mathbb{V}(\Delta(\tilde{T}_{s}))+2\mathrm{Cov}(H(\tilde{T}_{s}|X),\Delta(\tilde{T}_{s})).$
@@
$\mathbb{V}(\Delta(\tilde{T}_{s})\mid\tilde{T}_{s}\in\mathcal{T}_{\delta}^{s})\leq\frac{1}{4}L_{1}^{2}(b-a)^{2}.$
@@
$\mathrm{Cov}(H(\tilde{T}_{s}|X),\Delta(\tilde{T}_{s})\mid\tilde{T}_{s}\in\mathcal{T}_{\delta}^{s})\leq 0.$
@@
Combining these terms, we obtain a high-probability variance contraction:
$\mathbb{V}(H(\tilde{T}_{IB}|X)\mid\tilde{T}_{s}\in\mathcal{T}_{\delta}^{s})\leq\mathbb{V}(H(\tilde{T}_{s}|X)\mid\tilde{T}_{s}\in\mathcal{T}_{\delta}^{s})+\frac{1}{4}L_{1}^{2}(b-a)^{2}.$
Similarly, for $Y$,
$\mathbb{V}(H(\tilde{T}_{IB}|Y)\mid\tilde{T}_{s}\in\mathcal{T}_{\delta}^{s})\leq\mathbb{V}(H(\tilde{T}_{s}|Y)\mid\tilde{T}_{s}\in\mathcal{T}_{\delta}^{s})+\frac{1}{4}L_{1}^{2}(b-a)^{2},$
where $L_{1}=\prod_{i=1}^{d}\frac{1}{\lambda}$.
We can always choose $\lambda>1$ sufficiently large so that the Lipschitz constant $L_{1}=\prod_{i=1}^{d}\frac{1}{\lambda}$ is small enough to ensure $\frac{1}{4}L_{1}^{2}(b-a)^{2}+2\mathrm{Cov}(H(\tilde{T}_{s}|X),\Delta(\tilde{T}_{s})\mid\tilde{T}_{s}\in\mathcal{T}_{\delta}^{s})\leq 0$. Consequently, we obtain
$\mathbb{V}(H(\tilde{T}_{IB}|X)\mid\tilde{T}_{s}\in\mathcal{T}_{\delta}^{s})\leq\mathbb{V}(H(\tilde{T}_{s}|X)\mid\tilde{T}_{s}\in\mathcal{T}_{\delta}^{s}),$
@@
$\mathbb{V}(H(\tilde{T}_{IB}|Y)\mid\tilde{T}_{s}\in\mathcal{T}_{\delta}^{s})\leq\mathbb{V}(H(\tilde{T}_{s}|Y)\mid\tilde{T}_{s}\in\mathcal{T}_{\delta}^{s}).$

### v2

@@ C.5 Extended Results on ImageNet
Table 16: Accuracy (%) on CIFAR-10 with ResNet18. We report BatchNorm’s results since it significantly outperforms LayerNorm on CNNs.
Normalization Top-1 Top-5
BatchNorm 93.65 99.85
IterNorm 94.48 99.86
LayerNorm
RMSNorm
NormalNorm 94.01 99.86
IBNorm-S 95.49 99.85
IBNorm-L 95.57 99.87
IBNorm-T 95.51 99.83
C.6 The Effect of Compression Operation in IBNorm
We instantiate three representative forms of the compression function $f_{\lambda}$:
@@ C.6 The Effect of Compression Operation in IBNorm
These choices capture complementary compression behaviors. Specifically, IBNorm-S provides a linear baseline, IBNorm-L yields progressively stronger suppression of large activations, and IBNorm-T enforces a hard saturation effect on extreme values. Together, they span the spectrum from mild to strong tail suppression while sharing the desirable properties of boundedness and monotonicity. While other monotonic compression functions are possible, we provide these three forms to demonstrate the range of behaviors and provide a principled comparison.
@@
In conclusion, for these functionals, the compression ratio $\alpha_{\lambda}$ in Eqn. (7) is given by $\alpha_{\lambda}=1/\lambda$, which guarantees the bounded compression property whenever $\lambda\geq 1$. Compression operations in IBNorm compress the tail of activations while adjusting higher-order statistics.
@@
Distinction of IBNorm-S, IBNorm-L, and IBNorm-T IBNorm-S, IBNorm-L, and IBNorm-T are three types of IBNorm, which satisfy the compression property. Compression operations in IBNorm compress the tail of activations while adjusting higher-order statistics. IBNorm-L, IBNorm-T, and IBNorm-S exhibit different abilities to compress the tails of activations as shown in Figures 45, and 6. These differences in how aggressively each function compresses activation tails further lead to the distinction in the induced entropy of the output distribution. Because all three satisfy IB constraints, their IB curves (predictive MI, nuisance MI, token-level IB value) follow similar trends during training, making it difficult to distinguish their own characteristics purely from the IB framework. For this reason, we provide an additional quantitative analysis using kernel density entropy, which directly reflects how much information each compression function filters.
Specifically, to illustrate the distinct compression behaviors of IBNorm-S, IBNorm-L, and IBNorm-T, we compute the entropy of the kernel density estimate (KDE) of activations after applying each variant (with the same hyperparameter $\lambda=4$ under different input distributions. For an input Gaussian $\mathcal{N}(0,2)$ with entropy 1.768, the resulting KDE entropies are 1.695 for IBNorm-S, 1.530 for IBNorm-L, and 1.302 for IBNorm-T. Similarly, for an input Gaussian $\mathcal{N}(0,1)$ with entropy 1.418, the corresponding output entropies are 1.352 for IBNorm-S, 1.263 for IBNorm-L, and 1.192 for IBNorm-T. These results indicate that IBNorm-S performs the mildest compression, IBNorm-L achieves stronger compression, and IBNorm-T is the most aggressive in reducing entropy. This analysis highlights the distinct statistical effects of the three variants on latent representations.
Beyond compression strength, we observe differences in robustness. As shown in Tables 4 and 12, IBNorm-S is less sensitive to variations in the hyperparameter $\lambda$ compared to IBNorm-L and IBNorm-T. Similarly, ablation studies on the normalization structure shown in Tables 5 and 10 indicate that IBNorm-S is more robust to the ordering of the IB compression and standardization operations.
At the task level, performance differences align with these compression characteristics. Across diverse medium-scale LLMs (Llama 130M-1B and GPT 355M), IBNorm-S generally achieves the worst performance among the three IBNorm variants on evaluation tasks in Tables 1 and 6, reflecting its mildest compression. For challenging tasks, such as BBH and GPQA, the stronger compression applied by IBNorm-L can lead to improved performance, suggesting that mild compression can better support reasoning and generalization in difficult tasks.
C.7 Estimated Mutual Information Quantities and Token-level IB Value across Training
We track the evolution of three mutual information quantities during training: predictive information $\hat{I}(Y;T_{l})$, task-nuisance information $\hat{I}(T_{l-1};T_{l})$, and their token-level IB value $\hat{I}(Y;T_{l})-\hat{I}(T_{l-1};T_{l})$. These quantities are measured throughout training on test dataset for Llama-130M and GPT-2 (124M) on C4 and OpenWebText, respectively. The details in token-level IB value calculation are in Appendix G.
@@ C.7 Estimated Mutual Information Quantities and Token-level IB Value across Training
Figure 7: Evolution of predictive information $\hat{I}(Y;T_{l})$, task-nuisance information $\hat{I}(T_{l-1};T_{l})$, and the token-level IB value during training of Llama-130M on C4, using LN, IBNorm-T, and NormalNorm. Results are reported on the test set.
@@
Figure 8: Evolution of predictive information $\hat{I}(Y;T_{l})$, task-nuisance information $\hat{I}(T_{l-1};T_{l})$, and the token-level IB value during training of GPT-2 (124M) on OpenWebText, using LN, IBNorm-L, and NormalNorm. Results are reported on the test set.
@@
As shown in Figs. 78, and 9, these results demonstrate that IBNorm encourages representations that retain task-relevant information while suppressing irrelevant task-nuisance information, providing a explanation for the observed empirical gains.
C.8 Discussion on Training Time and Computational Overhead
IBNorm introduces only a single compression operation per normalization layer. Consequently, it does not incur significant computational overhead or cause numerical stability issues, even for larger models. Training time and memory usage of IBNorm remain comparable to standard variance-centric normalization methods such as LayerNorm or RMSNorm.
[…]
Table 17: Training time and VRAM usage for different normalization methods on LLaMA-60M.
Normalization Training Time ($\downarrow$) Total VRAM ($\downarrow$)
LayerNorm 1h 54min 95,944 MiB
RMSNorm 2h 15min 102,601 MiB
NormalNorm 3h 38min 139,207 MiB
IBNorm-S 2h 14min 102,460 MiB
IBNorm-L 2h 17min 106,371 MiB
IBNorm-T 2h 14min 104,853 MiB
[…]
Table 18: Training time and VRAM usage for different normalization methods on LLaMA-1B.
Normalization Training Time ($\downarrow$) Total VRAM ($\downarrow$)
LayerNorm 176h 02min 99,832 MiB
RMSNorm 190h 58min 110,574 MiB
NormalNorm 246h 06min 162,088 MiB
IBNorm-S 194h 29min 115,920 MiB
IBNorm-L 190h 02min 106,484 MiB
IBNorm-T 188h 41min 106,481 MiB
@@ Appendix D Details of Other Terms
In addition, we have the optimum-dependent term:
@@ E.1 Proof for Theorem 1
$\widehat{\text{IB}}_{S}(T_{IB})\geq\widehat{\text{IB}}_{S}(T_{s})\quad\text{almost surely}.$
@@
Given the dataset $S\sim\mathbb{P}_{X,Y}$ with size $M$ and one-layer neural networks, $f_{\text{1}}=\Phi_{\text{1}}\circ h$. Because $\eta$ is linear and invertible, it does not affect the mutual information terms (Hjelm et al., 2018; Ash and Doléans-Dade, 2000); therefore, it suffices to analyze the intermediate features $\tilde{T}_{s}:=\psi\circ\zeta\circ h(X)$ and $\tilde{T}_{IB}:=\psi\circ s_{\lambda}\circ\zeta\circ h(X)$. We analyze the IB value of intermediate features $\tilde{T}_{1}:=\psi\circ\zeta\circ h(X)$ in a general setting given a network $f$. First, we define the empirical gap $\Delta_{S}(\tilde{T}_{1};\tilde{T}_{1}^{*})$ between the IB bottleneck derived from arbitrary representation $\tilde{T}_{1}$ and the IB-optimum representation $\tilde{T}_{1}^{*}$, i.e.
@@
where the expectation and variance are taken with respect to the Lebesgue measure on $\mathcal{X}\times\mathcal{Y}$. The detailed proof is given in Appendix E.4.
Since $G(\cdot)$ is monotone in its variance arguments, Lemma 5 implies that, given arbitrary $\beta\in[0,1]$ and $\delta\in(0,1)$, with probability at least $1-\delta$ over the sample set $S\sim\mathbb{P}(X,Y)$ with size $M$, the following inequality holds:
@@
$\widehat{\text{IB}}_{S}(T_{IB})\geq\widehat{\text{IB}}_{S}(T_{s})\quad\text{almost surely}.$
@@ E.2 Proof for Lemma 3
$\Delta_{\text{S}}(\tilde{T}_{1};\tilde{T}^{*}_{1})\leq G(\sigma_{H_{1}|X},\sigma_{H_{1}|Y},\sigma_{p_{1}|X},\sigma_{p_{1}|Y},\mu_{p_{1}|X},\mu_{p_{1}|Y})+Q(T_{1}^{*}).$}
@@
Note that $\phi$ is a continuous, and concave function, and that $\lim\limits_{x\rightarrow 0}\phi(x)=0$. In addition, we have the following properties for the auxiliary function $\phi(\cdot)$ (Shamir et al., 2010):
@@
To bound the terms $A^{\prime}$ and $B^{\prime}$, we extend the results of Section 6.2 in (Shamir et al., 2010) to derive a general bound $|I(Y;T)-\hat{I}(Y;T)|$ under the setting where $T$ is a continuous random variable. Firstly, using the relationship I(Y;T)=H(T)H(T|Y)I(Y;T)=H(T)-H(T|Y), we apply the triangle inequality:
$|I(Y;T)-\hat{I}(Y;T)|\leq|H(T)-\hat{H}(T)|+|H(T|Y)-\hat{H}(T|Y)|$ (22)
Step 1: Bounding $|H(T)-\hat{H}(T)|$
By Lemma 6, for any two densities $p(t)$ and $\hat{p}(t)$, we have:
$|H(T)-\hat{H}(T)|=\left|\int_{\mathcal{T}}p(t)\log p(t)\,dt-\int_{\mathcal{T}}\hat{p}(t)\log\hat{p}(t)\,dt\right|\leq\int_{\mathcal{T}}\phi(|p(t)-\hat{p}(t)|)\,dt$ (23)
The difference in densities is:
$|p(t)-\hat{p}(t)|=\left|\sum_{x}p(t|x)(p(x)-\hat{p}(x))\right|$ (24)
Using the Cauchy-Schwarz inequality and the definition of $V(p(t|x))$ yields:
[…]
Substituting this back into the integral in Equation 23 gives:
[…]
Step 2: Bounding $|H(T|Y)-\hat{H}(T|Y)|$
We decompose the difference into two terms:
$|H(T|Y)-\hat{H}(T|Y)|\leq\left|\sum_{y}p(y)(H(T|y)-\hat{H}(T|y))\right|+\left|\sum_{y}(p(y)-\hat{p}(y))\hat{H}(T|y)\right|$ (27)
We have the following bound for the first term:
$\left|\sum_{y}p(y)(H(T|y)-\hat{H}(T|y))\right|$ $\leq\sum_{y}p(y)\int_{\mathcal{T}}\phi(|p(t|y)-\hat{p}(t|y)|)\,dt$ (28)
$\leq\sum_{y}p(y)\int_{\mathcal{T}}\phi\left(||\hat{p}(x|y)-p(x|y)||_{2}\sqrt{V(p(T=t|x))}\right)\,dt$ (29)
By the Cauchy-Schwarz inequality, the second term can be bounded as:
$\left|\sum_{y}(p(y)-\hat{p}(y))\hat{H}(T|y)\right|\leq||p(y)-\hat{p}(y)||_{2}\sqrt{V(\hat{H}(T|Y))}$ (30)
Step 3: Final Combination
Combining the bounds for $H(T)$ and $H(T|Y)$ in Equations 26 and 30 and extending the results in Section 6 (Shamir et al., 2010), the following inequality holds:
$|I(Y;T)-\hat{I}(Y;T)|\leq(2+\sqrt{2\log((|\mathcal{Y}|+2)/\delta)})\sqrt{\frac{V(\hat{H}(T|Y))}{M}}$ (31)
$\quad+2\int_{\mathcal{T}}\phi\Bigg((2+\sqrt{2\log((|\mathcal{Y}|+2)/\delta)})\sqrt{\frac{V(p(T|X))}{M}}\Bigg)p_{T}(t)\,dt$ (32)
$\leq(2+\sqrt{2\log((|\mathcal{Y}|+2)/\delta)})\sqrt{\frac{V({H}(T|Y))}{M}}$ (33)
$\quad+2\int_{\mathcal{T}}\phi\Bigg((2+\sqrt{2\log((|\mathcal{Y}|+2)/\delta)})\sqrt{\frac{V(p(T|X))}{M}}\Bigg)p_{T}(t)\,dt\quad(\text{Law of Large Numbers})$ (34)
Without loss of generality, Equation 31 implies:
@@
Thus, in our setting, we obtain:
@@
To bound the terms $A^{\prime}$ and $B^{\prime}$, we extend the results of Section 6.2 in (Shamir et al., 2010) to derive the corresponding bounds. Consequently, by applying the same bounding technique stated in I, we obtain:
@@
$|I(X;T)-\hat{I}(X;T)|\leq(2+\sqrt{2\log((|\mathcal{Y}|+2)/\delta)})\sqrt{\frac{V({H}(T|X))}{M}}$ (39)
@@
$|I(X;T^{*})-\hat{I}(X;T^{*})|\leq(2+\sqrt{2\log((|\mathcal{Y}|+2)/\delta)})\sqrt{\frac{V({H}(T^{*}|X))}{M}}$ (40)
@@
Thus, in our setting, we yield
@@
Combining I and II results in the following bound:
@@
where $Z(\tilde{T}_{0}^{*},\beta)$ is the partition function. Thus, we can write the remaining terms to optimum-dependent components:
@@ E.3 Proof for Proposition 4
$H(s_{\lambda}(Z))=H(Z)+\mathbb{E}[\log|\det J_{s_{\lambda}}(Z)|]\leq H(Z),$ (49)
because the Jacobian of the compression operator can be written as
$J_{s_{\lambda}}(Z)=D(Z)\!\left(I-\tfrac{1}{d}\mathbf{1}\mathbf{1}^{\top}\right),$ (50)
where $D(Z)=\mathrm{diag}\!\left(f_{\lambda}^{\prime}(|Z_{1}-\mu|),\ldots,f_{\lambda}^{\prime}(|Z_{d}-\mu|)\right)$. The projection matrix $I-\tfrac{1}{d}\mathbf{1}\mathbf{1}^{\top}$ has rank d1d-1 and operator norm at most $1$, and $f_{\lambda}^{\prime}(x)\leq\alpha_{\lambda}\leq 1$ by the bounded compression property. Consequently,
$\left|\det J_{s_{\lambda}}(Z)\right|\leq\alpha_{\lambda}^{\,d-1}\leq 1.$ (51)
@@ E.4 Proof for Lemma 5
$\tilde{T}_{s}=\Phi_{s}(X^{\prime})=\eta\circ\psi_{s}\circ\zeta(X^{\prime}),\qquad\tilde{T}_{IB}=\Phi_{IB}(X^{\prime})=\eta\circ\psi_{IB}\circ\zeta(X^{\prime}),$
where $X^{\prime}=h(X)$, $\psi_{IB}=\psi_{s}\circ s_{\lambda}$ with $s_{\lambda}$ being the compression operation in IBNorm.
@@
$H(\tilde{T}_{IB}|X)\leq H(\tilde{T}_{s}|X),\qquad H(\tilde{T}_{IB}|Y)\leq H(\tilde{T}_{s}|Y).$ (53)
If we consider the entropy $H$ in the discrete case, we can directly derive the relationship: consider the conditional entropy as a random variable over $X$ (or $Y$). Let $f^{\prime}(Z):=H(s_{\lambda}(Z))$, then the operator norm satisfies
$\|J_{s_{\lambda}}(z)\|_{\mathrm{op}}=\max_{1\leq i\leq d}\bigl|f_{\lambda}^{\prime}(|z_{i}-\mu|)\bigr|\leq\alpha_{\lambda}\leq 1,$ (54)
which implies that $s_{\lambda}$ is $\alpha_{\lambda}$-Lipschitz, where $\alpha_{\lambda}\leq 1$ by the bounded compression property of $f_{\lambda}$. Thus, $s_{\lambda}$ is a Lipschitz map with Lipschitz constant $L\leq 1$.
@@
In the differential entropy case, we cannot guarantee a global Lipschitz constant $L<1$ for $f^{\prime}(\tilde{T}_{s}|X)=H(\tilde{T}_{s}|X)+\Delta^{\prime}(\tilde{T}_{s}|X)$. Instead, we consider a high-probability setting for the continuous random variable $\tilde{T}_{IB}$ and $\tilde{T}_{s}$.
@@
$p_{\tilde{T}_{IB}|X=x}(t)=p_{\tilde{T}_{s}|X=x}(s_{\lambda}^{-1}(t))\,|\det J_{s_{\lambda}^{-1}}(t)|,$ (55)
@@
$H(\tilde{T}_{IB}|X=x)=H(\tilde{T}_{s}|X=x)+\Delta(\tilde{T}_{s}),\quad\Delta(\tilde{T}_{s}):=\log|\det J_{s_{\lambda}}(\tilde{T}_{s})|.$ (56)
@@
$\mathbb{V}(H(\tilde{T}_{IB}|X)\mid\tilde{T}_{s}\in\mathcal{T}_{\delta}^{s})=\mathbb{V}(H(\tilde{T}_{s}|X)+\Delta(\tilde{T}_{s})\mid\tilde{T}_{s}\in\mathcal{T}_{\delta}^{s}).$ (57)
Expanding Equation 57 using the variance formula gives
$\mathbb{V}(H(\tilde{T}_{s}|X)+\Delta(\tilde{T}_{s}))=\mathbb{V}(H(\tilde{T}_{s}|X))+\mathbb{V}(\Delta(\tilde{T}_{s}))+2\mathrm{Cov}(H(\tilde{T}_{s}|X),\Delta(\tilde{T}_{s})).$ (58)
@@
$\mathbb{V}(\Delta(\tilde{T}_{s})\mid\tilde{T}_{s}\in\mathcal{T}_{\delta}^{s})\leq\frac{1}{4}L_{1}^{2}(b-a)^{2}.$ (59)
@@
$\mathrm{Cov}(H(\tilde{T}_{s}|X),\Delta(\tilde{T}_{s})\mid\tilde{T}_{s}\in\mathcal{T}_{\delta}^{s})\leq 0.$ (60)
Combining Equations 59 and 59, we obtain:
$\mathbb{V}(H(\tilde{T}_{IB}|X)\mid\tilde{T}_{s}\in\mathcal{T}_{\delta}^{s})\leq\mathbb{V}(H(\tilde{T}_{s}|X)\mid\tilde{T}_{s}\in\mathcal{T}_{\delta}^{s})+\frac{1}{4}L_{1}^{2}(b-a)^{2}.$ (61)
Similarly, for $Y$,
$\mathbb{V}(H(\tilde{T}_{IB}|Y)\mid\tilde{T}_{s}\in\mathcal{T}_{\delta}^{s})\leq\mathbb{V}(H(\tilde{T}_{s}|Y)\mid\tilde{T}_{s}\in\mathcal{T}_{\delta}^{s})+\frac{1}{4}L_{1}^{2}(b-a)^{2},$ (62)
where $L_{1}=\prod_{i=1}^{d}\frac{1}{\lambda}$.
@@
Based on the bounded compression property and the property of $\Delta(\tilde{T}_{s})$, we have $\frac{1}{4}L_{1}^{2}(b-a)^{2}+2\mathrm{Cov}(H(\tilde{T}_{s}|X),\Delta(\tilde{T}_{s})\mid\tilde{T}_{s}\in\mathcal{T}_{\delta}^{s})\leq 0$. Thus, we obtain
$\mathbb{V}(H(\tilde{T}_{IB}|X)\mid\tilde{T}_{s}\in\mathcal{T}_{\delta}^{s})\leq\mathbb{V}(H(\tilde{T}_{s}|X)\mid\tilde{T}_{s}\in\mathcal{T}_{\delta}^{s}),$ (63)
@@
$\mathbb{V}(H(\tilde{T}_{IB}|Y)\mid\tilde{T}_{s}\in\mathcal{T}_{\delta}^{s})\leq\mathbb{V}(H(\tilde{T}_{s}|Y)\mid\tilde{T}_{s}\in\mathcal{T}_{\delta}^{s}).$ (64)
@@
$\mathbb{P}(\mathbb{V}(H(\tilde{T}_{IB}|X))\leq\mathbb{V}(H(\tilde{T}_{s}|X)))\geq 1-\delta,\quad\mathbb{P}(\mathbb{V}(H(\tilde{T}_{IB}|Y))\leq\mathbb{V}(H(\tilde{T}_{s}|Y)))\geq 1-\delta.$ (65)
@@
We have the following properties for the probability function $p(\cdot)$ and variance function $\mathbb{V}(\cdot)$ (Ash and Doléans-Dade, 2000):